Pls help solve dis equation, d sum of the ages of a father and mother is 81years.if the difference in the age is 9 years, calculate the age of the mother
Answer by ikleyn(52800)

.
Pls help solve dis equation, d sum of the ages of a father and mother is 81years.if the difference in the age is 9 years, calculate the age of the mother
--------------------------------------------

f + m = 81,   (1)
f - m =  9.   (2)

Add these two equations, and you easily find the father's age f = 45.
Then get the mother age m = 36.

If instead of (2) you have

m - f = 9    (2'),

then act correspondingly, and get m = 36, f = 45.

The condition doesn't say which of the two alternatives is really in the place.